### v3

@@ Abstract
In Estrada and Aros (2025), a relationship between gravitational tension (GT) and energy density, via the Kretschmann scalar (KS), was recently proposed to construct regular black hole (RBH) solutions in Pure Lovelock (PL) gravity. However, in PL gravity it is not viable to include a negative cosmological constant, as this leads to the appearance of a potential curvature singularity Cai and Ohta (2006). In this work, by choosing a particular set of coupling constants such that the resulting equations of motion for Lovelock gravity admit an $n$-fold degenerate ground state (LnFDGS) AdS solution, we construct an RBH solution with $\Lambda<0$, providing an energy density model analogous to (but distinct in its definition of GT) the one previously mentioned. Moreover, since relating the gravitational tension to the KS of the vacuum LnFDGS solution is nontrivial, we provide an alternative definition of both the KS and the GT. Remarkably, we obtain a model where there exists a value $r_{*}$ slightly greater than the extremal radius, $r_{*}>r_{ext}$, which could be on the order of the Planck length, such that the solutions of the vacuum AdS black hole and our AdS RBH become indistinguishable. However, at short length scales such that $r<r_{*}$, quantum effects would arise, causing both cases to differ in their geometry (suppressing the central singularity) and their thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior. We also provide a speculative methodology to compare theoretical results from Lovelock gravity with experimental results from the EHT for M87. Graphically, we have found that, depending on the chosen parameters, the theoretical results obtained for Lovelock gravity can match the experimental data.
@@ I Introduction
It is widely recognized that in the last decades various branches of theoretical physics predict the existence of extra dimensions. Although several experiments have attempted to test this hypothesis, it has not yet been observed. As a result, any theory that incorporates extra dimensions must be consistent with General Relativity in four dimensions. One such theory is Lovelock gravity. The Lagrangian of Lovelock gravity incorporates higher curvature terms as corrections to the Einstein-Hilbert action Lovelock (1971). Additionally, Lovelock’s theories respect the core principles of General Relativity; for instance, their equations of motion are of second order. It is worth noting that the specific case of Lovelock gravity known as Einstein-Gauss-Bonnet theory has gained attention in recent years for its applications in inflationary models and has been compared with the results from GW170817 Oikonomou (2021). See recent applications in references Bueno et al. (2025).
On the other hand, The singularity theorem, see Hawking and Ellis (2023); Penrose (1965), considered a triumph of theoretical physics, states in general terms that the collapse of matter could lead to the formation of singularities. However, it also suggests, broadly speaking, that the dynamics of gravity result in the breakdown of its own predictability. To address this fundamental issue, one could argue that the prediction of singularities arises from the incompleteness of current theories of gravity.
Although some theories that incorporate quantum effects, such as String Theory and Loop Quantum Gravity, to name just two examples, provide fundamentally different solutions to the singularity problem, it is not necessary to have a fully developed (fundamental) quantum gravity theory to offer solutions to the existence of singularities. This leads to ideas associated with the fact that gravity should be modified at small distances, typically around the Planck length. Therefore, in a sense, the removal or smoothing of these singularities should come from the incorporation of quantum effects in the matter fields present in the energy-momentum tensor. This latter idea has led to solutions known as Regular Black Holes (RBH), where the structure of the energy-momentum tensor is such that singularities at the origin are suppressed, generally leading to the formation of a core with a de Sitter structure rather than a singularity at the radial origin.
@@
One way that has attracted attention in recent years to define the energy-momentum tensor, in order to address the existence of singularities, is by establishing a relationship between the gravitational tension of the vacuum case (associated with the Kretschmann scalar) and the energy density. It is well known that an invariant associated with the measurement of tidal forces is the Kretschmann scalar Bena et al. (2021). In a four-dimensional, vacuum, spherically symmetric space, this scalar is proportional to $K\sim{M}^{2}/r^{6}$, leading to infinite tidal forces at the origin. Gravitational field tension in the spherically symmetric case is characterized by the curvature term given by the square root of the Kretschmann scalar of the vacuum solution, $F\sim\sqrt{K_{\text{Schw}}}\sim\frac{{M}}{r^{3}}$ in $4D$. This correlation is logical, as the spacetime tension should increase with the mass of the vacuum source Alencar et al. (2023). In this conection, as a way to address the problem of singularities while including an appropriate structure for the energy-momentum tensor, a higher-dimensional model was recently constructed in reference Estrada and Aros (2025) for Pure Lovelock (PL) gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar. Near the radial origin, where tidal forces and the gravitational tension in the vacuum solution diverge, the tidal forces in the model become finite. Specifically, the energy density is written as:
@@
being $F_{c},a$ constants. This model coincides with the RBH of Dymnikova Dymnikova (1992) in the four dimensional case. On the other hand, in a speculative sense, in four dimensions, this way of relating gravitational tension to energy density is analogous to the particle production ratio in the quantum Schwinger effect DYMNIKOVA (1996). We can find studies related to this latter topic in the references Ansoldi (2008); Alencar et al. (2023); Estrada and Muniz (2023) . However, this is a topic that requires deeper investigation.
On the other hand, over the last 50 years, much has been discussed regarding how General Relativity and other theories of gravity can be improved to address theoretical inconsistencies and mismatches with observations. Naturally, in recent decades, there has been a growing interest in determining the role of the cosmological constant in the functioning of gravity. Although the expansion of the universe has usually been associated with the existence of a positive cosmological constant, this fact brings some complications from a theoretical standpoint. Firstly, the presence of a positive cosmological constant is associated with the existence of a cosmological horizon, which prevents the existence of an asymptotic regime, making the unambiguous identification of the ADM mass of a black hole problematic. Secondly, a positive cosmological constant is usually associated with negative pressure, implying thermodynamic instability.
On the contrary, the presence of a negative cosmological constant offers good explanations for some physical ambiguities. A spacetime with a negative cosmological constant allows for the definition of an asymptotic boundary in the geometry of Penrose diagrams. In this context, the existence of an asymptotic boundary enables the correct definition of conserved charges Aros et al. (2001), such as mass/energy. The existence of a well-defined boundary in an AdS universe has led to the development of the AdS/CFT theory. In broad terms, the AdS/CFT theory is a correspondence between a quantum theory defined in a certain class of anti-de Sitter space and a conformal field theory defined on its boundary with one lower dimension. Moreover, it is worth mentioning that in recent years, attention has been drawn to associating the thermodynamics of an AdS Schwarzschild space, where the negative cosmological constant has been linked to a positive thermodynamic pressure Kubiznak and Mann (2012). This has allowed for the study of phase transitions, where stable regions can be found. In this way, it is of physical interest to study how the presence of a negative cosmological constant influences the thermodynamic properties of black holes.
@@
A special case of Lovelock theories is Pure Lovelock theory (PL) Cai and Ohta (2006); Dadhich et al. (2012). The latter theory mentioned considers only a single term in the Lagrangian. Related to the aforementioned, a disadvantage of PL theory lies in the fact that considering the presence of a negative cosmological constant in a spherically symmetric spacetime leads to the appearance of a curvature singularity. This singularity is such that, although the metric tensor is regular for a radius greater than the event horizon, the Ricci and Kretschmann invariants diverge at a point within this location. This issue has not been extensively investigated, and thus, it may be related to a potential breakdown of spacetime. Therefore, studying the recent analogy between gravitational tension and energy density, as mentioned earlier, seems to be unfeasible in PL theory under the presence of a negative cosmological constant for the construction of RBH solutions.
@@
One potential drawback of generic Lovelock gravity is the existence of multiple ground states, meaning more than one solution with constant curvature spaces, or equivalently, more than one potential effective cosmological constant Camanho and Edelstein (2013). These effective cosmological constants can be complex numbers, which can lead to instability of the ground states under dynamical evolution. To address this issue, one can choose the coupling constant in such a way that the equations of motion roughly take the form $(R-\Lambda)^{n}=0$ Crisostomo et al. (2000). In this case, there is a single ground state, but it is n-fold degenerate with constant curvature. This scenario is referred to as Lovelock with a unique vacuum (LUV) or Lovelock with an n-fold degenerate ground state (LnFDGS). In the context of this work, this coupling constant structure allows for the observation of a single negative effective cosmological constant, which, among other things, has facilitated the development of techniques for determining conserved charges Aros and Estrada (2019). This also seems to make sense in the framework of the AdS/CFT correspondence, as two or more negative cosmological constants would lead to different asymptotic boundary structures for the same background. See some recent applications of LnFDGS theory in references Xu and Yung (2019); Aros and Estrada (2019); Estrada and Aros (2019); Aros et al. (2022); Estrada (2022); Aros and Estrada (2024)
@@
PL has significant differences compared to the LnFDGS theory Aros and Estrada (2019); Crisostomo et al. (2000). While PL considers only a single term from the Lovelock series, LnFDGS includes all the terms from this series up to $n=[d/2]$, where the square brackets represent the floor function. Moreover, PL cannot be factorized into several equivalent negative effective cosmological constants like LnFDGS. Remarkably, LnFDGS does not exhibit a curvature singularity of the type mentioned earlier. The latter offers the advantage of displaying a well-defined asymptotically AdS structure. In the context of this work, the absence of the aforementioned pathology in LnFDGS allows for the construction of RBH solutions with the presence of a negative cosmological constant, and for testing the physical consequences of its presence from both the geometric and thermodynamic perspectives.
@@
On the other hand, it is well known that light is not directly observable at the event horizon of a black hole. Instead, what can be seen is the so-called shadow of a black hole, a dark region surrounded by light that results from the gravitational lensing phenomenon around the event horizon. In this way, an important achievement of the Event Horizon Telescope (EHT) collaboration has been the construction of images of the shadows of the supermassive black holes M87 Akiyama et al. (2021) and Sgr A Akiyama et al. (2022). In this regard, testing the behavior of shadows in scenarios that consider extra dimensions has also attracted attention: see, for example, reference Nozari et al. (2024) for Scalar-Tensor-Vector theories and reference Singh et al. (2023) (and the references therein) for rotating black hole shadows. It is important to mention that in reference Paithankar and Kolekar (2023), connections were established between the parameters of the black hole solution and the structure of the shadow for extra-dimensional and asymptotically flat scenarios. This study was carried out for scenarios where the event horizon radius can be determined analytically. Thus, relationships were established between the size of the latter and the radii of the photon sphere and the shadow. The vacuum solutions of Pure Lovelock and $f(R)$ gravity were studied. In this way, it becomes interesting to establish numerical relationships between the aforementioned shadow characteristics for Lovelock theories, which lack an analytical expression for their event horizon and are not asymptotically flat, but instead possess an AdS asymptote.
@@
Regarding the aforementioned points, in this work, we will provide a methodology to study the analogy between gravitational tension and energy density in order to construct regular black hole solutions in LnFDGS gravity under the presence of a negative cosmological constant. It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution in the same way as for PL theory is complicated, because several additional terms would appear in the gravitational tension. Due to this, unlike Estrada and Aros (2025), we will define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory, and consequently, a redefinition of the gravitational tension. Additionally, we will speculatively discuss the connection between our model and the quantum Schwinger effect for this theory. We will test how the presence of the negative cosmological constant influences the obtained RBH solution, both in terms of its geometric and thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we will propose a method to numerically and graphically obtain the aforementioned relationships and analyze their physical behavior. We also provide a speculative way to compare theoretical results obtained for Lovelock gravity with the experimental results from the EHT for M87.
@@
II A brief revision of Lovelock gravity with $n$ fold degenerated ground state

### v4

@@ Abstract
In Estrada and Aros (2025), a link between gravitational tension (GT) and energy density via the Kretschmann scalar (KS) was proposed to construct regular black holes (RBHs) in Pure Lovelock (PL) gravity. However, including a negative cosmological constant in PL gravity leads to a curvature singularity Cai and Ohta (2006). Here, we choose the coupling constants such that the Lovelock equations admit an $n$-fold degenerate AdS vacuum (LnFDGS), allowing us to construct an RBH with $\Lambda<0$, where the energy density is analogous to the previously mentioned model. To achieve this, we propose alternative definitions for both the KS and GT. We find that, for mass parameter values greater than the extremal value $M_{\text{min}}$, our RBH solution becomes indistinguishable from the AdS vacuum black hole from inside the event horizon out to infinity. At small scales, quantum effects modify the geometry and thermodynamics, removing the singularity. Furthermore, due to the lack of analytical relationships between the event horizon, photon sphere, and shadow in LnFDGS, we propose a numerical method to represent these quantities.
@@ I Introduction
The detection of gravitational waves Abbott et al. (2016, 2017a) has significantly strengthened the position of General Relativity (GR), even surpassing its already well-established success at the Solar System scale Damour (2007). Nevertheless, some inconsistencies of GR at quantum scales, along with the still unknown nature of dark energy and dark matter, have driven the scientific community to investigate modified theories of gravity. As a first example of these alternative frameworks, the Horndeski scalar-tensor theory Horndeski (1974) generalizes GR by introducing one or more scalar degrees of freedom. This theory is carefully constructed so that its equations of motion prevent the emergence of an unwanted extra degree of freedom, commonly referred to as a ’ghost’,which would carry negative energy. As a second example, we consider the $f(R)$ theories of gravity Nojiri and Odintsov (2006). As noted in reference Nojiri et al. (2009), this framework offers a gravitational alternative for a unified description of inflation, dark energy, and dark matter, without the need to manually introduce an inflaton field or additional dark components. It is also worth mentioning that black hole solutions have been studied within this theory de la Cruz-Dombriz et al. (2009).
@@
In line with the previous paragraph, in recent decades, modified gravity theories predicting the existence of extra dimensions have also attracted significant attention. Although several experiments have attempted to test this hypothesis, it has not yet been observed. As a result, any theory that incorporates extra dimensions must be consistent with General Relativity in four dimensions. One such theory is Lovelock gravity. The Lagrangian of Lovelock gravity incorporates higher curvature terms as corrections to the Einstein-Hilbert action Lovelock (1971). Additionally, Lovelock’s theories respect the core principles of General Relativity; for instance, their equations of motion are of second order. It is worth noting that one particular case of Lovelock theories, namely, Einstein-Gauss-Bonnet gravity, has gained attention in recent years in the context of inflationary models and has been confronted with the GW170817 observational results Oikonomou (2021); Odintsov and Oikonomou (2020).
On the other hand, it is well known that vacuum solutions describing black holes possess a physical singularity, referring to a location where curvature invariants diverge and the tidal forces near a black hole’s singularity become infinite, which can lead to the unbounded stretching of an object, a phenomenon known as spaghettification. One way to address the existence of physical singularities is to define appropriate forms for the matter sources in the energy–momentum tensor such that the central singularity is suppressed. Black holes that are free of a physical singularity are usually referred to as regular black holes (RBHs).
@@
In recent years, the relationship between the gravitational tension associated with the Kretschmann scalar in the vacuum case and the energy density has been investigated to eliminate the existence of singularities. It is well known that an invariant associated with the measurement of tidal forces is the Kretschmann scalar Bena et al. (2021). In a four-dimensional, vacuum, spherically symmetric space, this scalar is proportional to $K\sim{M}^{2}/r^{6}$, leading to infinite tidal forces at the origin. Gravitational field tension in the spherically symmetric case is characterized by the curvature term given by the square root of the Kretschmann scalar of the vacuum solution, $F\sim\sqrt{K_{\text{Schw}}}\sim\frac{{M}}{r^{3}}$ in $4D$. This correlation is logical, as the spacetime tension should increase with the mass of the vacuum source Alencar et al. (2023). In this conection, as a way to address the problem of singularities while including an appropriate structure for the energy-momentum tensor, a higher-dimensional model was recently constructed in reference Estrada and Aros (2025) for Pure Lovelock (PL) gravity, where the energy density encodes the gravitational information of the vacuum solution through the Kretschmann scalar. Near the radial origin, where tidal forces and the gravitational tension in the vacuum solution diverge, the tidal forces in the model become finite. Specifically, the energy density is written as:
@@
In summary, the model proposed in Estrada and Aros (2025) is such that its energy density encodes the gravitational information of the vacuum solution. Thus, near the central singularity, where the gravitational tension and tidal forces of the vacuum black hole become infinite, the energy density of the model remains finite, suppressing the singularity and giving rise to a regular black hole (RBH). This model coincides with the RBH of Dymnikova Dymnikova (1992) in the four-dimensional case. On the other hand, in a speculative sense, in four dimensions, this way of relating gravitational tension to energy density is analogous to the particle production ratio in the quantum Schwinger effect DYMNIKOVA (1996). We can find studies related to this latter topic in the references Ansoldi (2008); Alencar et al. (2023); Estrada and Muniz (2023) . However, this is a topic that requires deeper investigation.
Since the origins of general relativity, there has been an interest in understanding the role of the cosmological constant. The observed accelerated expansion of the universe can be attributed to a small, yet strictly positive, four dimensional cosmological constant of approximately $\Lambda\sim 10^{-121}\mbox{\,Planck units}\sim 10^{-52}[m^{-2}]\sim 10^{-3}[(\textrm{eV})^{4}]$ Barrow and Shaw (2011). Unfortunately, a positive cosmological constant introduces several technical complications. First, the presence of a positive cosmological constant leads to the existence of a cosmological horizon; thus, it prevents the definition of a proper asymptotic regime and complicates an unambiguous identification of a black hole’s ADM mass. Second, a positive cosmological constant is commonly linked to negative pressure, which implies thermodynamic instability.
During the last three decades, primarily due to the AdS/CFT conjecture, considering a negative cosmological constant has been of paramount interest in theoretical physics. From a theoretical point of view, considering asymptotically AdS spaces has advantages, such as the construction of a proper action principle —one that admits the definition of an asymptotic boundary where well-defined boundary conditions can be imposed and where unambiguous conserved charges can be computed.
Despite the above mentioned issues regarding physical problems addressed by including a negative cosmological constant, there has been little speculation in the literature about a possible numerical negative value for this constant. For example, references Zakharov (2014, 2018), using observations of stars near the Galactic Center, estimate a negative value for the cosmological constant: in Zakharov (2014), for anti–de Sitter spacetimes corresponding to a negative $\Lambda$ term of approximately $0.4\times 10^{-20}\text{ m}^{-2}$. The authors claim that this value is consistent with some ideas within the framework of a certain class of multidimensional string models. In Zakharov (2018), also using observations of stars near the Galactic Center, the orbital precession in the Reissner–Nordström–AdS solution is studied, obtaining values of $4.68\times 10^{-39}\text{ cm}^{-2}<\Lambda<0$. The authors state that their constraints on the cosmological constant are much weaker not only than their cosmological estimates but also than the constraints obtained from Solar System data.
In the context of this work, that is, within Lovelock theories, some authors assume that the coupling constants of the theory could be of the order of Planck scale quantities Chernicoff et al. (2016). This makes sense if one considers that the theory might only be testable at high energies. As we will see in detail below, the zeroth-order coupling constant is proportional to the cosmological constant. Under this idea, a negative cosmological constant would be of the order of powers of the Planck length, $\Lambda\sim-\ell_{p}^{-2n}$. That is, an AdS radius on the order of the Planck length. However, there are also works that define intermediate energies as those which are weak enough to be testable at energy scales below the Planck scale, but still high enough to feel the effects of the higher-curvature terms in Lovelock theory Camanho et al. (2016). Under this latter assumption, $\Lambda\sim-\bar{\ell}^{-2n}$, with $\bar{\ell}>\ell_{p}$. That is, an AdS radius sufficiently larger than the Planck length, in order to satisfy what has been mentioned above.
When higher dimensions ($d>4$) are considered, several families of gravitational theories emerged that satisfy the same conditions of General Relativity, such as second-order differential equations. Arguably, the simplest of these families is the $d$-dimensional Lovelock gravity, which corresponds to the addition of all of the lower-dimensional topological densities. General relativity corresponds to the simplest case.
@@
A special case of Lovelock theories is Pure Lovelock theory (PL) Cai and Ohta (2006); Dadhich et al. (2012). The latter theory mentioned considers only a single term in the Lagrangian. Related to those discussed above, a disadvantage of PL theory is that considering the presence of a negative cosmological constant in a spherically symmetric spacetime leads to the appearance of a curvature singularity. This singularity is such that, although the metric tensor is regular for a radius greater than the event horizon, the Ricci and Kretschmann invariants diverge at a point within this location. This issue has not been extensively investigated; therefore, it may be related to a potential breakdown of spacetime. Thus, studying the recent analogy between gravitational tension and energy density, as carried out in Estrada and Aros (2025), appears to be unfeasible within the framework of PL theory in the presence of a negative cosmological constant for the construction of RBH solutions.
@@
To address the issue above—namely, the apparent impossibility of studying the recent analogy between gravitational tension and energy density in pure Lovelock theory under the influence of a negative cosmological constant—this work will take the following into account: One potential drawback of generic Lovelock gravity is the existence of multiple ground states, meaning more than one solution with constant curvature spaces, or equivalently, more than one potential effective cosmological constant Camanho and Edelstein (2013). These effective cosmological constants can be complex numbers, which can lead to instability of the ground states under dynamical evolution. To address this issue, one can choose the coupling constant in such a way that the equations of motion roughly take the form $(R-\Lambda)^{n}=0$ Crisostomo et al. (2000). In this case, there is a single ground state, but it is n-fold degenerate with constant curvature. This scenario is referred to as Lovelock with an n-fold degenerate ground state (LnFDGS). This is also known as the degenerate vacuum. In the context of this work, this coupling constant structure allows for the observation of a single negative effective cosmological constant, which, among other things, has facilitated the development of techniques for determining conserved charges Aros and Estrada (2019). This also seems to make sense in the framework of the AdS/CFT correspondence, as two or more negative cosmological constants would lead to different asymptotic boundary structures for the same background. See some recent applications of LnFDGS theory in references Xu and Yung (2019); Aros and Estrada (2019); Estrada and Aros (2019); Aros et al. (2022); Estrada (2022); Aros and Estrada (2024)
@@
PL has significant differences compared to the LnFDGS theory Aros and Estrada (2019); Crisostomo et al. (2000). While PL considers only a single term from the Lovelock series, LnFDGS includes all the terms from this series up to $n=[d/2]$, where the square brackets represent the floor function. Moreover, PL cannot be factored into several equivalent negative effective cosmological constants, such as LnFDGS. Remarkably, LnFDGS does not exhibit a curvature singularity of the type mentioned earlier. The latter offers the advantage of displaying a well-defined asymptotically AdS structure. In the context of this work, the absence of the aforementioned pathology in LnFDGS enables the construction of RBH solutions with a negative cosmological constant, allowing for the testing of the physical consequences of its presence from both geometric and thermodynamic perspectives.
@@
On the other hand, it is well known that light is not directly observable at the event horizon of a black hole. Instead, what can be detected is the so-called shadow of a black hole, a dark region surrounded by light, resulting from gravitational lensing around the event horizon. One of the major achievements of the Event Horizon Telescope (EHT) collaboration has been the imaging of the shadows of the supermassive black holes M87 Akiyama et al. (2021) and Sgr A Akiyama et al. (2022). It is worth mentioning that, given the theoretical interest driven by the presence of a negative cosmological constant, the study of shadows in an asymptotically AdS regime has also been explored in some references (see, for example, Belhaj et al. (2021); Hamil et al. (2024)). In this context, there has also been strong motivation, from a theoretical perspective, to study shadows in extra-dimensional scenarios. See, for example, reference Nozari et al. (2024) for Scalar-Tensor-Vector theories, and reference Singh et al. (2023) (and references therein) for rotating black hole shadows. See also Vagnozzi and Visinelli (2019). As previously mentioned, the study of the physical properties of black holes in Lovelock gravity has attracted considerable attention in recent years Aros and Estrada (2019); Xu and Yung (2019); Estrada and Aros (2019); Aros et al. (2022); Estrada (2022); Aros et al. (2001); Bueno et al. (2025). Along this line, it is natural to investigate the physical properties of their shadows. It is worth noting that in reference Paithankar and Kolekar (2023), analytical relations were established for the photon sphere radius and the shadow radius of singular vacuum black holes in pure Lovelock (PL) gravity. However, for most specific cases within Lovelock theory, such analytical results are not attainable. Therefore, it is of physical interest to develop a numerical methodology for representing black hole shadows.
@@
Regarding the points above, in this work, we will provide a methodology to study the analogy between gravitational tension and energy density, aiming to construct regular black hole solutions in LnFDGS gravity in the presence of a negative cosmological constant. It is worth mentioning that directly relating the gravitational tension to the Kretschmann scalar of the vacuum LnFDGS solution, in the same way as for PL theory, is complicated because several additional terms would appear in the gravitational tension. Due to this, unlike Estrada and Aros (2025), we will define an alternative version of the Kretschmann scalar suitable for the LnFDGS AdS theory, and consequently, a redefinition of the gravitational tension. We will test how the presence of the negative cosmological constant influences the obtained RBH solution, both in terms of its geometric and thermodynamic properties. Additionally, since it is not possible to find analytical relationships between the event horizon, the photon sphere radius, and the shadow size in LnFDGS, we propose a method to numerically and graphically obtain these relationships and analyze their physical behavior.
@@
II A brief revision of Lovelock gravity with n-fold degenerated ground state